# SPECIAL RELATIVITY

### v9

@@ 2.9 Mistake in Commonly Used Method of Coupling Fields and Particles
However, there is a common mistake made in accelerator physics connected with the difference between $\vec{x}(t)$ and $\vec{x}_{cov}(t)$ trajectories. Let us look at this difference from the point of view of electrodynamics of relativistically moving charges. To evaluate fields arising from external sources we need to know their velocity and positions as a function of the lab frame time $t$. Suppose one wants to calculate properties of radiation. Given our previous discussion the question arises, whether one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particle moving along non-covariant trajectories like $\vec{x}(t)$. We claim that the answer to this question is negative. This algorithm for solving usual Maxwell’s equations in the lab frame, which is considered in all standard treatments as relativistically correct, is at odds with the theory of special relativity. This essential point has never received attention in the physical community. Only the solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle trajectories in the lab frame.
2.10 Clarification of the True Content of the ”Single Frame” Theory
Let us now examine the logical content of the concept of a ”single inertial frame”. If a traveler in a co moving frame, similar to an observer in the lab frame, introduces a definite coordinate-time grid, there is always a definite transformation between these two four-dimensional coordinate systems. Thus, particle trajectories are always viewed from the lab frame as a result of successive transformations, and the form of these transformations depends on the choice of coordinate systems in the comoving and the lab frame.
One might well wonder why it is necessary to discuss how different inertial frames are related to one another. The point is that all natural phenomena follow the principle of relativity, which is a restrictive principle: it says that the laws of nature are the same (or take the same form) in all inertial frames. In agreement with this principle, usual Maxwell’s equations can always be exploited in any inertial frame where electromagnetic sources are at rest using Einstein synchronization procedure in the rest frame of the source. The fact that one can deduce electromagnetic field equations for arbitrary moving sources by studying the form taken by Maxwell’s equations under the transformation between rest frame of the source and the frame where the source is moving is a practical application of the principle of relativity. The question now arises how to assign a time coordinate to the lab frame.
Coordinates serve the purpose of labeling events in an unambiguous way, and this can be done in infinitely many different ways. The principle of relativity dictates that Maxwell’s equations can be applied in the lab frame only in the case when Lorentz coordinates are assigned and particle trajectories are viewing from the lab frame as a result of successive infinitesimal Lorentz transformations between the lab and comoving inertial frames.
A ”single frame” (non-covariant 3+1) approach to relativistic particle dynamics has been used in particle tracking calculations for about seventy years. However, the type of clock synchronization which provides the time coordinate $t$ in the corrected Newton’s equation has never been discussed in literature. It is clear that without an answer to the question about the method of synchronization used, not only the concept of velocity, but also the dynamics law has no physical meaning. A ”single frame” approach to relativistic particle dynamics is forcefully based on a definite synchronization assumption but this is actually a hidden assumption. According to conventional particle tracking, the dynamical evolution in the lab frame is based on the use of the lab frame time $t$ as an independent variable. Such approach to relativistic particle dynamics is actually based on the use of a not standard (not Einstein) clock synchronization assumption in the lab frame.
In fact, the usual for accelerator engineering study of relativistic particle motion in a constant magnetic field looks precisely the same as in nonrelativistic Newtonian mechanics and the trajectories of the electrons does not include relativistic kinematics effects. According to textbooks, this is no problem. If no more than one frame is involved, one does not need to use (and does not need to know) the theory of relativity. Only when one passes from one reference frame to another the relativistic context is important. Conventional particle tracking in a constant magnetic field is actually based on classical Newton mechanics. It is generally believed that the electrodynamics problem, similar to conventional particle tracking, can be treated within a description involving a single inertial frame and one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particles moving along the non-covariant (single frame) trajectories.
This is misconception. The situation when only one frame is involved and the relativistic context is unimportant cannot be realized. The lab observer may argue, ”I don’t care about other frames.” Perhaps the lab observer doesn’t, but nature knows that, according to the principle of relativity, Maxwell’s equations are always valid in the Lorentz comoving frame. Electrodynamics equations can be written down in the lab frame only when a space-time coordinate system has been specified. An observer in the lab frame has only one freedom. This is the choice of a coordinate system (i.e. the choice of clock synchronization convention) in the lab frame. After this, the theory of relativity states that the electrodynamics equations in the lab frame are the result of transformation of Maxwell’s equations from the Lorentz comoving frame to the lab frame.
2.11 Bibliography and Notes
1. Many physicists tend to think of Galilean transformations as pre-relativistic transformations between spatial coordinates and time that are not compatible with the special theory of relativity. To quote e.g. Bohm [1] ”… the Galilean law of addition of velocities implies that the speed of light should vary with the speed of the observing equipment. Since this predicted variation is contrary to the fact, the Galilean transformations evidently cannot be the correct one.”. Similar statements can also be found in recently published pedagogical papers. To quote e.g. Drake and Purvis [2] ”One of the great insights to come relativity theory was the realization that Galilean transformations are wrong. The correct way to translate the space-time measure of events between inertial frames is with the Lorentz transformations” However, this is not true. Galilean transformations are simply transformations relating a given coordinate set to another coordinate set. The space-time continuum can be described in arbitrary coordinates, and choice of this set of coordinates cannot change the geometry of space-time.
2. The mathematical argument that in the process of transition to arbitrary coordinates the geometry of the space-time does not change, is considered in textbooks as erroneous. To quote L. Landau and E. Lifshitz [3]: ”This formula is called the Galileo transformation. It is easily to verify that this transformation, as was to be expected, does not satisfy the requirements of the theory of relativity; it does not leave the interval between events invariant.”. This fact is ascribed to a lack of understanding of the difference between convention-dependent and convention-invariant parts of the theory. In pseudo-Euclidean geometry the interval between events is an invariant in arbitrary coordinates. A comparison with three-dimensional Euclidean space might help here. In the usual 3D Euclidean space, one can consider a Cartesian coordinate system ($x,y,z)$, a cylindrical coordinate system ($r,\phi,z$), a spherical coordinate system ($\rho,\theta,\phi$), or any other. Depending on the choice of the coordinate system one respectively has $ds^{2}=dx^{2}+dy^{2}+dz^{2}$, $ds^{2}=dr^{2}+r^{2}d\phi^{2}+dz^{2}$, $ds^{2}=d\rho^{2}+\rho^{2}d\theta^{2}+\rho^{2}\sin\theta^{2}d\phi^{2}$. The metric actually does not change, but the components of the metric do, depending only on the choice of coordinates. In general, in fact, we write $ds^{2}=g_{ik}dx^{i}dx^{k}$. Considering Cartesian coordinates we will always have $g_{ij}=\mathrm{diag}(1,1,1)$. Similarly, Lorentz transformations between inertial frames with Einstein coordinates leave the components of the metric tensor unvaried.
3. There is a widespread belief that a Lorentz transformation reduces to a Galilean transformation in the non-relativistic limit. To quote e.g. French [4]: ”The reduction of $t^{\prime}=\gamma(t-vx/c^{2})$ to Galilean relation $t^{\prime}=t$ requires $x\ll ct$ as well as $v/c\ll 1$”. Similar statements can also be found in recently published textbooks. To quote Rafelski [5]: ”A wealth of daily experience shows that the Galilean coordinate transformation (GT) is correct in the nonrelativistic limit in which the speed of light is so large that it plays no physical role. Any coordinate transformation replacing the GT must also agree with this experience, and thus must contain the GT in the nonrelativistic limit.” This statement is absurd conclusion even from a mathematical standpoint. This has been recognized by some expects, perhaps most explicitly by Baierlein who states [6] that ”If the Lorentz transformation for infinitesimal $v/c$ were to reduce to the Galilean transformation, then iterative process could never generate a finite Lorentz transformation that is radically different from the Galilean transformation. But the finite transformations are radically different, and so -however subtly-the infinitesimal Lorentz transformation must differ significantly from the Galilean transformation.”
4. Many authors of textbooks still attribute a measurable status to the conventional quantities. To quote e.g. Cristodoulides [7] ”The fact that Galilean transformation does not leave Maxwell’s equations has already been mentioned […] On the other hand, experiments show that the speed of light in vacuum is independent of the source or observer.”.
5. Because we have empirical access only to the round-trip average speed of light, statements about the magnitude and isotrophy of the one-way speed of light must reflect the assumptions made in the choice of time coordinatization, and such entities change as the theory is re-synchronized. To quote C. Anderson, I. Vetharaniam, and G. Stedman [8]: ”No experiment, then, is a ”one-way” experiment. An empirical test of any property of the one-way speed of light is not possible. Such quantities as the one-way speed of light are irreducibly conventional in nature, and recognizing this aspect is to recognize a profound feature of nature”.
6. The peculiarity of the kinematic consequence of using Galilean transformations is that the speed of light emitted by a moving source depends on the relative velocity between source and observer. A widespread theoretical argument used to support the incorrectness of Galilean transformations is the conclusion that a Galilean transformation of the velocity of light is not consistent with the explanation of reflection and refraction. This idea is a part of the material in well-known books. To quote Pauli [9] ”[…] it is essential that the spherical waves emitted by the dipoles in the body should interfere with the incident wave. If we now think of the body as at rest, and the light source moving relative to it, then […] the wave emitted by the dipoles will have velocity different from that of the incident wave. Interference is therefore not possible.” This conclusion is incorrect. It is clear that the incident and scattered wave at any given point have the same frequency and can interfere.
7. For a general discussion of the difference between path and trajectory we suggest reading the book [10].
8. The standard textbooks erroneously identify the properties of space-time with the familiar form that certain synchronization-dependent quantities assume under the Einstein’s clock synchronization. These quantities are, for example, the formulae for ”time dilation” and ”length contraction”. Fortunately, it has also been occasionally stressed in the literature that the forms of the above ”relativistic effects” are coordinate-dependent, their forms depend in turn on the kind of synchronization procedure adopted. To quote Leubner, K. Auflinger, and P. Krumm [11]: ”… there is a widespread belief among students that the familiar form of coordinate-dependent quantities like the measured velocity of light, the Lorentz transformation between two observers, ”addition of velocities”, ”time dilation”, ”length contraction” which they assume under the standard clock synchronization, is relativity proper. In order to demonstrate that this is by no means so, this paper studies the consequences of a non-standard synchronization, and it is shown that drastic changes in the appearance of all these quantities are thus induced.”
9. Clarification of the true content of the non-covariant theory can be found in various advanced textbooks. To quote e.g. Ferrarese and Bini [12]: ”… within a single inertial frame, the time is an absolute quantity in special relativity also. As a consequence, if no more than one frame is involved, one would not expect differences between classical and relativistic kinematics. But in the relativistic context there are differences in the transformation laws of the various relative quantities (of kinematics or dynamics), when passing from one reference frame to another.” We see that authors give a special role to concept of a ”single inertial frame”. The name ”single inertial frame” tends to suggest that a distinctive trait of non-covariant theory is the absence of relativistic kinematics in the description of particle motion. This point was expressed by Friedman [15]: ”Within any single inertial frame, things looks precisely the same as in Newtonian kinematics: there is an enduring Euclidean three-space, a global (i.e. absolute) time $t$, and law of motion. But different inertial frames are related to one another in a non-Newtonian fashion.” According to conventional particle tracking, within the ”single” frame there is no relativistic kinematics effects. This, as we already mentioned, contradicts the Maxwell’s electrodynamics.
@@ 3 Space-Time and Its Coordinatization
3.1 Introductory Remarks
Let us discuss an ”operational interpretation” of the Lorentz and absolute time coordinatizations. We should underline that we claim the non covariant approach to relativistic particle dynamics is actually based on the use of a not standard and unusual clock synchronization assumption within the theory of relativity. It is important to know how to operationally interpret the absolute time convention i.e. how one should perform the clock synchronization in the lab frame. The result is very interesting, since it tell us about difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view.
@@ 3.2 Choice of Coordinates System in an Inertial Frame
Each physical phenomenon occurs in space and time. A concrete method for representing space and time is a frame of reference. One-and-the same space and time can be represented by various coordinate-time grids, i.e., by various frames of references. Even the simplest space-time coordinate systems require carefully description.
Clocks reveal the motion of a particle through the coordinate-time grid. The general approach to the determination of the motion of a particle is the following: at any instant a particle has a well-defined velocity $\vec{v}$ as measured in a laboratory frame of reference. How is a velocity of a particle found? The velocity is determined once the coordinates in the lab frame are chosen, and is then measured at appropriate time intervals along the particle’s trajectory. But how to measure a time interval between events occurring at different points in space? In order to do so, and hence measure the velocity of a particle within a single inertial lab frame, one first has to synchronize distant clocks. The concept of synchronization is a key concept in the understanding of special relativity. It is possible to think of various methods to synchronize distant clocks (1). The choice of a convention on clock synchronization is nothing more than a definite choice of coordinates system in an inertial frame of reference of the Minkowski space-time.
The space-time continuum can be described in arbitrary coordinates. By changing these arbitrary coordinates, the geometry of the four-dimensional space-time obviously does not change, and in the special theory of relativity we are not limited in any way in the choice of a coordinates system. Relying on the geometric structure of Minkowski space-time, one defines the class of inertial frames and adopts a Lorentz frame with orthonormal basis vectors. Within the chosen Lorentz frame, Einstein’s synchronization procedure of distant clocks (which based on the constancy of the speed of light in all inertial framers) and Cartesian space coordinates are enforced.
@@ 3.3 Inertial Frame where a Source of Light is at Rest
Let us give an ”operational interpretation” of the Lorentz coordinatizations. The fundamental laws of electrodynamics are expressed by Maxwell’s equations, according to which, as well-known, light propagates with the same velocity $c$ in all directions. This is because Maxwell’s theory has no intrinsic anisotropy. It has been stated that in their original form Maxwell’s equations are only valid in inertial frames. However, Maxwell’s equations can be written down in coordinate representation only if the space-time coordinate system has already been specified.
@@
The problem of assigning Lorentz coordinates to the lab frame in the case of accelerated motion is complicated. We would like to start with the simpler question of how to assign space-time coordinates to an inertial frame, where a source of light is at rest. We need to give a ”practical”, ”operational” answer to this question. The most natural method of synchronization consists in putting all the ideal clocks together at the same point in space, where they can be synchronized. Then, they can be transported slowly to their original places (slow clock transport) (2).
@@
The usual Maxwell’s equations are valid in any inertial frame where sources are at rest and the procedure of slow clock transport is used to assign values to the time coordinate. The same considerations apply when charged particles are moving in non-relativistic manner. In particular, when oscillating, charged particles emit radiation, and in the non-relativistic case, when charges oscillate with velocities much smaller than $c$, dipole radiation is generated and described with the help of the Maxwell’s equations in their usual form.
@@
Let’s examine in a more detail how the dipole radiation term comes about. The retardation time in the integrands of the expression for the radiation field amplitude, can be neglected in the cases where the trajectory of the charge changes little during this time. It is easy to find the conditions for satisfying this requirement. Let us denote by $a$ the order of magnitude of the dimensions of the system. Then the retardation time $\sim a/c$. In order to ensure that the distribution of the charges in the system does not undergo a significant change during this time, it is necessary that $a\ll\lambda$, where $\lambda$ is the radiation wavelength. Thus, the dimensions of the system must be small compared to radiation wavelength. This condition can be written in still another form $v\ll c$, where $v$ is of the order of magnitude of the velocities of the charges. In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory. Therefore, one should not be surprised to find that dipole radiation theory gives fields very much like the instantaneous theory.
@@
The theory of relativity offers an alternative procedure of clocks synchronization based on the constancy of the speed of light in all inertial frames. This is usually considered a postulate but, as we have seen, it is just a convention. The synchronization procedure that follows is the usual Einstein synchronization procedure. Suppose we have a dipole radiation source. When the dipole light source is at rest, the field equations are constituted by the usual Maxwell’s equations. Indeed, in dipole radiation theory we consider the small expansion parameter $v/c\ll 1$ neglecting terms of order $v/c$. In other words, in dipole radiation theory we use zero order non relativistic approximation. Einstein synchronization is defined in terms of light signals emitted by the dipole source at rest, assuming that light propagate with the same velocity $c$ in all directions. Using Einstein synchronization procedure in the rest frame of the dipole source, we actually select the Lorentz coordinate system.
@@
Slow transport synchronization is equivalent to Einstein synchronization in inertial system where the dipole light source is at rest (3). In other words, suppose we have two sets of synchronized clocks spaced along the $x$ axis. Suppose that one set of clocks is synchronized by using the slow clock transport procedure and the other by light signals. If we would ride together with any clock in either set, we could see that it has the same time as the adjacent clocks, with which its reading is compared. This is because in our case of interest, when light source is at rest, field equations are the usual Maxwell’s equations and Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions. Using any of these synchronization procedures in the rest frame we actually select a Lorentz coordinate system. In this coordinate system the metric of the light source has Minkowski form $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$. We want now to turn to the subject of electromagnetic waves. In free space, the electric field $\vec{E}$ of an electromagnetic wave emitted by the dipole source at rest satisfies the equation $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$.

### v10

@@ 2.9 Mistake in Commonly Used Method of Coupling Fields and Particles
However, there is a common mistake made in accelerator physics connected with the difference between $\vec{x}(t)$ and $\vec{x}_{cov}(t)$ trajectories. Let us look at this difference from the point of view of electrodynamics of relativistically moving charges. To evaluate fields arising from external sources we need to know their velocity and positions as a function of the lab frame time $t$. Suppose one wants to calculate the properties of radiation. Given our previous discussion, the question arises, whether one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by a particle moving along non-covariant trajectories like $\vec{x}(t)$. We claim that the answer to this question is negative. This algorithm for solving Maxwell’s equations in the lab frame, which is considered in all standard treatments as relativistically correct, is at odds with the theory of special relativity. This essential point has never received attention in the physical community. Only the solution of the dynamics equations in covariant form gives the correct coupling between Maxwell’s equations and particle trajectories in the lab frame.
@@ 3.1 Choice of Coordinates System in an Inertial Frame
Let us discuss an ”operational interpretation” of the Lorentz and absolute time coordinatizations. Each physical phenomenon occurs in space and time. A concrete method for representing space and time is a frame of reference. One-and-the-same space and time can be represented by various coordinate-time grids, i.e., by various frames of reference. Even the simplest space-time coordinate systems require careful description.
Clocks reveal the motion of a particle through the coordinate-time grid. The general approach to the determination of the motion of a particle is the following: at any instant, a particle has a well-defined velocity $\vec{v}$ as measured in a laboratory frame of reference. How is the velocity of a particle found? The velocity is determined once the coordinates in the lab frame are chosen and is then measured at appropriate time intervals along the particle’s trajectory. But how to measure a time interval between events occurring at different points in space? In order to do so, and hence measure the velocity of a particle within a single inertial lab frame, one first has to synchronize distant clocks. The concept of synchronization is a key concept in the understanding of special relativity. It is possible to think of various methods to synchronize distant clocks  [^10] According to the thesis of the conventionality of simultaneity [13, 14, 15, 16, 17], the simultaneity of distant events is a conventional matter, as it can be legitimately fixed in different manners in any given inertial reference frame. To quote e.g. Moeller [13]: ”All methods for the regulation of clocks meet with the same fundamental difficulty. The concept of simultaneity between two events in different places has no exact objective meaning at all since we cannot specify any experimental method by which this simultaneity could be ascertained. The same is therefore true also for the concept of velocity.”. The choice of a convention on clock synchronization is nothing more than a definite choice of coordinates system in an inertial frame of reference of the Minkowski space-time.
The space-time continuum can be described in arbitrary coordinates. By changing these arbitrary coordinates, the geometry of the four-dimensional space-time does not change, and in the special theory of relativity, we are not limited in any way in the choice of a coordinates system. Relying on the geometric structure of Minkowski space-time, one defines the class of inertial frames and adopts a Lorentz frame with orthonormal basis vectors. Within the chosen Lorentz frame, Einstein’s synchronization procedure of distant clocks (which is based on the constancy of the speed of light in all inertial frames) and Cartesian space coordinates are enforced.
@@ 3.2 Inertial Frame where a Source of Light is at Rest
Let us give an ”operational interpretation” of the Lorentz coordinatizations. The fundamental laws of electrodynamics are expressed by Maxwell’s equations, according to which, as well-known, light propagates with the same velocity $c$ in all directions. This is because Maxwell’s theory has no intrinsic anisotropy. It has been stated that in their original form, Maxwell’s equations are only valid in inertial frames. However, Maxwell’s equations can be written down in coordinate representation only if the space-time coordinate system has already been specified.
@@
The problem of assigning Lorentz coordinates to the lab frame in the case of light source motion is complicated. We would like to start with the simpler question of how to assign space-time coordinates to an inertial frame, where a source of light is at rest. We need to give a ”practical”, and ”operational” answer to this question. The most natural method of synchronization consists in putting all the ideal clocks together at the same point in space, where they can be synchronized. Then, they can be transported slowly to their original places (slow clock transport)  [^11] In the text first published in 1923, Eddington discussed, apparently for the first time, a procedure for synchronization using slow transport of clocks [18]. The details can be found in review [8] ( see also [19]). .
@@
The usual Maxwell’s equations are valid in any inertial frame where sources are at rest and the procedure of slow clock transport is used to assign values to the time coordinate. The same considerations apply when charged particles are moving in a non-relativistic manner. In particular, when oscillating, charged particles emit radiation, and in the non-relativistic case, when charges oscillate with velocities much smaller than $c$, dipole radiation is generated and described with the help of Maxwell’s equations in their usual form.
@@
Let’s examine in more detail how the dipole radiation term comes about. The retardation time in the integrands of the expression for the radiation field amplitude can be neglected in the cases where the trajectory of the charge changes little during this time. It is easy to find the conditions for satisfying this requirement. Let us denote by $a$ the order of magnitude of the dimensions of the system. Then the retardation time $\sim a/c$. To ensure that the distribution of the charges in the system does not undergo a significant change during this time, it is necessary that $a\ll\lambda$, where $\lambda$ is the radiation wavelength. Thus, the dimensions of the system must be small compared to the radiation wavelength. This condition can be written in still another form $v\ll c$, where $v$ is of the order of magnitude of the velocities of the charges. In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory. Therefore, one should not be surprised to find that dipole radiation theory gives fields very much like the instantaneous theory.
@@
The theory of relativity offers an alternative procedure of clock synchronization based on the constancy of the speed of light in all inertial frames. This is usually considered as a postulate but, as we have seen, it is just a convention. The synchronization procedure that follows is the usual Einstein synchronization procedure. Suppose we have a dipole radiation source. When the dipole light source is at rest, the field equations are constituted by the usual Maxwell’s equations. Indeed, in dipole radiation theory we consider the small expansion parameter $v/c\ll 1$ neglecting terms of order $v/c$. In other words, in dipole radiation theory we use zero-order nonrelativistic approximation. Einstein synchronization is defined in terms of light signals emitted by the dipole source at rest, assuming that light propagates with the same velocity $c$ in all directions. Using the Einstein synchronization procedure in the rest frame of the dipole source, we select the Lorentz coordinate system.
@@
Slow transport synchronization is equivalent to Einstein’s synchronization in an inertial system where the dipole light source is at rest  [^12] Many authors of textbooks still attribute a reality status of the one-way speed of light. To quote Hrasko [20]: ”It is sometimes claimed that Einstein’s synchronization of distant clocks $A$ and $B$ is circular. … This argument is, however, fallacious. It is true that one-way measurement of light velocity can indeed be performed only if clocks at the endpoints are synchronized correctly. But since they need not show the correct coordinate time, they can be synchronized without light signals by transporting them from the common site in a symmetrical manner. …. As we see, the thought experiment described is capable of proving the constancy of light speed if it is true or to disprove it if it is false. It provided, therefore, solid logical foundation for Einstein’s synchronization prescription.” This logical argument is incorrect. Slow clock transport synchronization is equivalent to Einstein’s synchronization only in an inertial system where the light source is at rest.. In other words, suppose we have two sets of synchronized clocks spaced along the $x$ axis. Suppose that one set of clocks is synchronized by using the slow clock transport procedure and the other by light signals. If we would ride together with any clock in either set, we could see that it has the same time as the adjacent clocks, with which its reading is compared. This is because in our case of interest when the light source is at rest, field equations are the usual Maxwell’s equations and Einstein synchronization is defined in terms of light signals emitted by a source at rest assuming that light propagates with the same velocity $c$ in all directions. Using any of these synchronization procedures in the rest frame we select a Lorentz coordinate system. In this coordinate system, the metric of the light source has Minkowski form $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$. We want now to turn to the subject of electromagnetic waves. In free space, the electric field $\vec{E}$ of an electromagnetic wave emitted by the dipole source at rest satisfies the equation $\Box^{2}\vec{E}=\nabla^{2}\vec{E}-\partial^{2}\vec{E}/\partial(ct)^{2}=0$.